OpenStudy (elonasushchik):

Medal and Fan!!! Line AB contains points A (8, -4) and B (1, -5). The slope of line AB is

OpenStudy (elonasushchik):

-7

OpenStudy (elonasushchik):

-1 7

OpenStudy (elonasushchik):

1 7

OpenStudy (elonasushchik):

7

OpenStudy (elonasushchik):

@Bee_see and @SyedMohammed98

syed98 (syedmohammed98):

use the same formula i:e m= y-y/(x-x) so it's m =-5-(-4) / 1 - 8 m = -1/7.

syed98 (syedmohammed98):

umm i showed the solution above...

OpenStudy (elonasushchik):

@Agl202

OpenStudy (elonasushchik):

k. so its B. tx! @SyedMohammed98

OpenStudy (elonasushchik):

tx!!

syed98 (syedmohammed98):

You're Welcome

OpenStudy (agl202):

=)

OpenStudy (mathstudent55):

A (8, -4) and B (1, -5) \(m = \dfrac{-5 - (-4)}{1 - 8} \) \(m = \dfrac{-5 + 4}{-7} \) \(m = \dfrac{-1}{-7} \) \(m = \dfrac{1}{7} \) Answer is choice C.